A triode has mutual conductance = 2.0 millimho and plate resistance = 20 kΩ. It is desired to amplify a signal by a factor of 30. What load resistance should be added in the circuit?

Grade:upto college level

1 Answers

Aditi Chauhan
9 years ago
Sol. r base p = 20 KΩ = Plate Resistance Mutual conductance = g base m = 2.0 milli mho = 2 * 10^–3 mho Amplification factor μ = 30 Load Resistance = RL = ? We know A = μ/1 + r base p/R base L where A = voltage amplification factor ⇒ A = r base p * g base m/1 + r base p/R base L where μ = r base p * g base m ⇒ 30 = 20 *10^3 *2 * 10^-3/1 + 20000/R base L ⇒ 3 = 4R base L/R base L + 20000 ⇒ 3R base L + 60000 = 4 R base L ⇒ R base L = 60000 Ω = 60 KΩ